# MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations

### v1

@@ Abstract
Large language models (LLMs) have become pervasive in our everyday life. Yet, a fundamental obstacle prevents their use in many critical applications: their propensity to generate fluent, human-quality content that is not grounded in reality. The detection of such hallucinations is thus of the highest importance. In this work, we propose a new method to flag hallucinated content, MMD-Flagger. It relies on Maximum Mean Discrepancy (MMD), a non-parametric distance between distributions. On a high-level perspective, MMD-Flagger tracks the MMD between the generated documents and documents generated with various temperature parameters. We show empirically that inspecting the shape of this trajectory is sufficient to detect most hallucinations. This novel method is benchmarked on two machine translation datasets, on which it outperforms natural competitors.
@@ 1 Introduction
Neural probabilistic modeling (bengio_et_al_2003, ) in an auto-regressive framework has become the de facto paradigm for natural language processing. Powered by the transformer architecture (Vaswani-2017-attention, ), such neural networks have achieved impressive performances in text translation (kocmi_et_al_2024, ), abstractive text summarization (shakil_farooq_kalita_2024, ), generative question-answering (daull_et_al_2023, ), and many other applications. More recently, the trend has been to bring forward so-called Large Language Models (LLMs) such as GPT (openai_2024, ), which can tackle simultaneously nearly all these different tasks. Popularized by ChatGPT (a chatbot version of GPT), they are used daily by dozens of millions of users (Yuyang-2023-BigData, ).
However, LLMs have a key limitation which prevents their use in critical applications: their propensity to generate hallucinations. Commonly characterized by the generation of fluent but factually inaccurate or unsubstantiated content, these are observed in all models (ji_et_al_2023, ), although rare (Guerreiro-2023-looking, ). Within the domain of machine translation, the focus of this work, this phenomenon manifests as translations that, while grammatically correct, deviate semantically from the contextual information present in the source language (Guerreiro-2023-looking, ). We provide an example in Table 1. Hallucinated content is a critical drawback, especially in domains where information accuracy is essential, such as law Benkirane-2024-machine . In these fields, getting wrong (yet seemingly correct) information even once can undermine users’ trust permanently. Detecting hallucinations at test time remains a challenging task.
Focusing on standard encoder-decoder architecture, it is believed that the decoding step is one of the main causes of hallucinations (ji_et_al_2023, ). In practice, beam search is commonly used to generate the output token sequence, given its superior performance, even with respect to maximum a posteriori (Meister-etal-2020-beam, ). This is particularly true for machine translation tasks (Guerreiro-2023-looking, ; Xu-etal-2023-understanding, ). Nevertheless, beam search tends to suffer from the so-called likelihood trap (das-balke-2022-quantifying, ): a high-likelihood token sequence often results in a low-quality token sequence.
An alternative approach is simply to sample the next token according to the probability distribution given by the decoder’s softmax layer at each step. The main advantage in doing so is the stochasticity of outputs, giving rise to many diverse output documents from a single input. A further refinement is to introduce a temperature parameter ($\tau$ in our notation), which controls the randomness of the next token prediction; a low value (say $<0.1$) corresponding roughly to a probability distribution concentrated on the most likely token, and a higher value (say $>1$) making the distribution near uniform. Thus, it is not surprising that, in practice, low temperatures are reported to induce low variability in the sampled token sequences (Peeperkorn-etal-2024, ), while high temperatures result increase in the hallucination rate (Renze-2024-effect, ; Huang-2025, ).
@@
Table 1: Examples of hallucinated translation from the LFAN-HALL dataset generated by beam search ($y_{\text{beam}}$), and stochastic samples with varying temperature parameters ($y_{\text{sto}}^{\tau}$). The MMD distances between $y_{\text{beam}}$ and $y_{\text{sto}}^{\tau=0.5}$ as a function of $\tau$ can be visualized as the red line in Figure 1.
@@
$y_{\rm beam}$ The staff were very friendly and helpful. The room was clean and comfortable.
@@
The central objective of this study is to compare the default output of a language model (generated using beam search) with stochastic samples obtained under varying temperature. At low temperature settings, we observe that the sampled outputs are highly similar to the default output if the model is not hallucinating, and moderately dissimilar if it is. We believe this occurs because of the following reason: when a clear optimal candidate token exists, both sampling (at low temperatures) and beam search tend to favor the most probable token. However, in cases where no single candidate clearly dominates (often indicative of hallucination), the outputs are more likely to diverge. In contrast, at high temperature, generation becomes increasingly random, and sampled outputs tend to be substantially different from the default output regardless of the model’s behavior. This pattern is illustrated in Table 1. As a consequence, plotting the similarity between the default output and the stochastic outputs yields two distinct curves: either a monotonously increasing trajectory in the non-hallucinating scenario, or a $U$-shape curve in the hallucinating scenario. This can be visualized in Figure 1.
Drawing on these findings, we introduce a novel hallucination detection method: MMD-Flagger. The key idea of the method is to simply i) plot the similarity between default output and stochastic samples as a function of the temperature, and ii) automatically detect whether this curve is monotonously increasing or $U$-shaped. In the latter case, we flag the output as hallucination. As a measure of similarity, we propose to use the Maximum Mean Discrepancy (MMD) (Gretton-2012a, ). As a shape detection mechanism, we simply estimate the inflection point of the trajectory and check whether it is located near 00 (see Figure 1).
The paper is organized as follows. In Section 2, we detail the operating procedure of MMD-Flagger after recalling the necessary background on temperature sampling and MMD. We benchmark our method against several natural competitors in Section 3, before concluding in Section 5. The codebase is available at https://github.com/Kensuke-Mitsuzawa/MMD-Flagger.
Figure 1: MMD-Flagger at work on the LFAN-HALL dataset: two outputs, generating two MMD trajectories. In red, the trajectory for the hallucinated translation (see Table 1), in green, the trajectory corresponding to the correct one. Text boxes show examples of $y_{\rm beam}$ and samples from $Y_{\rm sto}^{\tau}$ (two of 25 samples), generated by stochastic decoding with temperatures $\tau$. We flag $y_{\rm beam}$ as a hallucination when the minimum value of $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ (the vertical black dashed line) falls within the red region. Otherwise, the output is deemed plausible (minimum falling into the green region).
@@ 1.1 Related work
Hallucination detection is a very active field, with a wide range of approaches proposed in recent years. Let us review the most related work; We refer to (ji_et_al_2023, ) for a more complete survey of mitigation methods in various domains.
Uncertainty quantification. Following the taxonomy outlined by (Huang-2025, ), MMD-Flagger falls under the umbrella of uncertainty estimation, which aims to quantify a model’s confidence in its output. Stemming from the observation that models are more confident when they do not hallucinate, the Monte-Carlo dropout-based MC-Dsim (Fomicheva-2020-unsupervised, ; Zerva-2021-unbabel, ) estimates the stability of generated token sequences. A similar approach is to inspect the log-probability of output sequences (Guerreiro-2023-looking, ) and to flag as hallucination if this log-probability is too low (Seq-Logprob). Alternatively, quality estimation methods (Rei-etal-2022-cometkiwi, ; Feng-etal-2022-language, ) correlate low-quality translations with hallucinations, though they are not originally designed for hallucination detection. Empirical evidence also suggests that combining multiple detectors improves detection accuracy (Guerreiro-2023-looking, ; Himmi-2024-enhanced, ). In the context of QA, (Varshney-2023stitchtimesavesnine, ) proposes to use the minimum token probability as an indicator of uncertainty, based on the premise that lower token probabilities suggest higher uncertainty. We compare MMD-Flagger to MC-Dsim and Seq-Logprob in Section 3.
External knowledge. Beyond uncertainty estimation, several approaches aim to detect hallucination by leveraging external knowledge. For example, (Min-etal-2023-factscore, ) introduce factuality scores derived from knowledge graphs. In machine translation, (Xu-etal-2023-understanding, ) propose a supervised hallucination detection model trained on synthetic pseudo-hallucination datasets. The LMvLM framework (Cohen-etal-2023-lm, ) employs an examiner language model to interrogate an examinee model through multi-turn interactions, identifying potential inconsistencies. Similarly, the Chain-of-Verification method (dhuliawala-etal-2024-chain, ) makes the model’s reasoning process explicit, enabling humans to assess the validity of its outputs. We do not compare against such methods since MMD-Flagger does not use external information.
Model internals. Building on the assumption that models “know when they hallucinate” (duan_et_al_2024, ), other approaches look at models’ internal states to detect hallucinations. (Ferrando-etal-2022-towards, ; Dale-2022-detecting, ) measure how much the input contributes to the output and show that one can detect hallucination when the source sentence contribution is low. It is also possible to analyze input-output relationships from an approach using optimal transport (Guerreiro-etal-2023-optimal, ). We showcase in Appendix G how MMD-Flagger can be adapted to leverage information from the model internals. However, the more consistent experimental results are achieved using only the model’s outputs.
Kernel methods and LLMs. Another key ingredient of our approach is the use of MMD to compute similarities between outputs of LLMs. To the best of our knowledge, only one prior study connects MMD with LLMs. (Zhangs-2024MMDMP, ) propose a two-sample testing framework that employs MMD as the test statistic to distinguish between human-written and machine-generated texts. Thus, their goal is not hallucination detection, but rather authenticity classification, making the application of MMD to hallucination detection a novel contribution of our work. More similar to our approach, (Nikitin-2024-kernel, ) proposes kernel language entropy, a kernel-based method to quantify the uncertainty in LLM outcomes, and hallucination detection is proposed based on the entropy score (Farquhar-2024detecting, ).
@@ 2 Methodology
In this section, we describe the proposed method, MMD-Flagger. Given an input document $x$, the starting point of the method is to generate the output $y=f(x)$ to be inspected, using beam search. The next step is to generate additional documents by stochastic sampling at various temperature parameters, which is detailed in Section 2.1. Then, the MMD between $y$ and the newly created groups is computed for each temperature parameter. We detail this computation in Section 2.2. At this point, for a given input / output pair $(x,y)$, we have obtained a series of MMD values corresponding to different temperature settings. We explain our procedure to inspect this trajectory in Section 2.3.
2.1 Stochastic sampling and temperature
In this section, we present the encoder-decoder architecture considered in this work and detail the role of the temperature in stochastic sampling.
Encoder-decoder architecture.
We consider the encoder-decoder architecture (Sutskever-2014, ) with self-attention mechanism (Vaswani-2017-attention, ). Given an input token sequence $x$, the encoder $e$, which consists in multiple layers of self-attention and feed-forward networks encodes the sequence into a representation with fixed length $T$. The decoder then generates the output sequence autoregressively, conditioned on both the encoded fixed-length vector and the already processed tokens. More precisely, each output token is generated as
$Y_{t}=\Phi(Y_{t-1},\ldots,Y_{1},e(x))\,,$
where $\Phi$ is a (potentially) stochastic sampler. This process iterates until the EOS token is sampled. In our framework, the way sampling operates is based on the assumption that the decoder is equipped with a softmax layer converting logits scores $\mathbb{R}^{T\times V}$ into a discrete probability distribution $P\in\mathbb{R}^{V}$, where $V$ is the size of the dictionary.
Argmax. Arguably, the most natural thing to do is to select the token associated with the highest $P_{i}$ as output. We refer to this as argmax. Simply picking the most likely token comes with a lot of rigidity, since some tokens can attract a lot of mass, such as the EOS token. Alternatively, it is very common to sample a random token using this distribution. There are many ways to do so.
@@ 2.1 Stochastic sampling and temperature
Stochastic sampling. A first possibility is simply to sample according to $P$. That is, sample a random variable with distribution given by $P$ and select the associated token. We denote $Y_{\text{sto}}=\text{LLM}_{\text{sto}}(x)$ the random output obtained by stochastic sampling.
Temperature parameter. A variation on the previous theme is to introduce a temperature parameter $\tau>0$. The softmax transformation is modified into
$P_{i}=\frac{\mathrm{exp}(z_{i}/\tau)}{\sum_{j=1}^{V}\mathrm{exp}(z_{j}/\tau)}\,,$ (1)
@@
where $z_{i}$ represents the logit score for the $i$-th token. As recalled in Section 1, high temperature gives a lot of variability (Yehuda-etal-2024, ) (equivalent to uniform sampling in the limit), while low temperature degrades to argmax sampling.
Beam Search. Rather than considering only the most likely next token, beam search allows for a broader search over possible sequences by keeping multiple candidates. This principle underlies beam search, a deterministic decoding algorithm widely used in the literature (Meister-etal-2020-beam, ). Formally,
$\begin{split}S_{t}&=\operatorname{argmax}_{S^{\prime}\subseteq\mathcal{B}_{t},|S^{\prime}|=w}\log p_{\theta}(S^{\prime}|x),\\ \text{where}\qquad\mathcal{B}_{t}&=\{s_{t-1}\circ s\mid s\in\mathcal{V}\text{ and }s_{t-1}\in S_{t-1}\}\,.\end{split}$
In the previous display, $S$ is the stack of candidate token sequences maintained during beam search, $S_{t}$ is the stack at the $t$-th time step, $\mathcal{B}_{t}$ is the candidate set of next possible token sequences, and $\mathcal{V}$ indicates the vocabulary list, and $s_{t-1}\circ s$ is the operation of adding a sampled token $s$. Beam search usually starts with $S_{0}=\{\text{BOS}\}$, and holds $w$-candidates of token sequences in $S$. Beam search can keep the $w$-best path of tokens in selecting the $t$-th token, while the argmax sampling simply chooses the token with the highest probability value. We denote by $y_{\text{beam}}=\text{LLM}_{\text{beam}}(x)$ the token sequence generated by beam search. Note that it is also possible to use beam-search with a temperature parameter, but we do not do so in this work, as is typical. While aiming to find the most likely token sequence, beam search is known to sometimes generate low-quality outputs having a high likelihood (das-balke-2022-quantifying, ). Stochastic sampling offers a potential mitigation for this issue by introducing diversity into the generation process. We aim to compare the sequences generated by beam search with those produced through stochastic sampling. For a temperature parameter $\tau\in T$, we collect a set of token sequences $Y^{\tau}_{\rm sto}=\{\text{LLM}^{\tau}_{\text{sto}}(x)\}_{n=1}^{N}$ generated by stochastic sampling. To summarize, in this work we compare $y_{\text{beam}}$ and $Y^{\tau}_{\rm sto}$ for various values of $\tau$.
Alternative sampling strategies. To conclude this section, we note that there exist many other alternative sampling strategies, such as top-k (fan-etal-2018-hierarchical, ) or nucleus sampling (Holtzman-2020, ). They can mitigate the hallucinations by placing constraints on sampling the next tokens (Huang-2025, ).

### v3

@@ Abstract
Large language models (LLMs) have become pervasive in our everyday life. Yet, a fundamental obstacle prevents their use in many critical applications: their propensity to generate fluent, human-quality content that is not grounded in reality. The detection of such hallucinations is thus of the highest importance. In this work, we propose a new method to flag hallucinated content: MMD-Flagger. It relies on Maximum Mean Discrepancy (MMD), a non-parametric distance between distributions. On a high-level perspective, MMD-Flagger tracks the MMD between the output to inspect and counterparts generated with various temperature parameters. We show empirically that inspecting the shape of this trajectory is sufficient to detect most hallucinations. This novel method is benchmarked on machine translation and summarization datasets, on which it exhibits competitive performance relative to natural competitors.
MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations
Kensuke Mitsuzawathanks: Corresponding author. Université Côte d’Azur, CNRS, LJAD, France kensuke.mitsuzawa@unice.fr          Damien Garreau Center for Artificial Intelligence and Data Science (CAIDAS) Julius-Maximilans-Universität Würzburg, Germany damien.garreau@uni-wuerzburg.de
@@ 1 Introduction
Neural probabilistic modelling in an auto-regressive framework (Bengio et al., 2003), powered by Transformers (Vaswani et al., 2017), underpins modern NLP, enabling advances in translation (Kocmi et al., 2024), summarization (Shakil et al., 2024), question answering (Daull et al., 2023), and more. Large Language Models (LLMs) such as GPT (Radford et al., 2018), popularized by ChatGPT and now used by millions (Deng et al., 2023), can address diverse tasks such as the machine translation or summarization. But it can suffer from hallucinations, fluent yet factually incorrect generated content (Ji et al., 2023). Though relatively infrequent (Guerreiro et al., 2023b), hallucinations are especially concerning in translation and summarization, where outputs may deviate semantically from the source (see Table 1), and in high-stakes domains such as law (Benkirane et al., 2024), where even a single error can erode trust. Given that detecting hallucinations at the test time remains a major challenge (Sriramanan et al., 2024), we focus on detection in white-box settings and inspecting a single LLM’s output. This allows us to retain full internal access to the model’s parameters and intermediate states.
The emergence of hallucinations in generated output is involved with the decoding strategy used for next-token prediction (Shi et al., 2024). While no universally optimal strategy exists, beam search (Meister et al., 2020) and its variants are often practical choices, particularly for models fine-tuned through alignment with human feedback (Shi et al., 2024). However, beam search and its variants are prone to the so-called likelihood trap (Das and Balke, 2022), where sequences with high probability can nevertheless yield poor-quality or uninformative outputs, which can cause the hallucination.
An alternative decoding strategy is to sample the next token directly from the probability distribution given by the model’s softmax layer. The key advantage of this approach is its stochasticity, which allows diverse outputs to be generated from the same input. A common refinement introduces a temperature parameter ($\tau$), which adjusts the sharpness of the probability distribution. Stochastic sampling, therefore, provides opportunities to improve generation quality under appropriate parameter choices (Shi et al., 2024).
@@
Table 1: Examples of hallucinated translations from the LFAN-HALL dataset generated by beam search ($y_{\text{hyp}}$), and stochastic samples with varying temperature parameters ($y_{\text{sto}}^{\tau}$). The MMD distances between $y_{\text{hyp}}$ and $y_{\text{sto}}^{\tau=0.5}$ as a function of $\tau$ can be visualized as the “Hallucinated” in Figure 1.
@@
$y_{\rm hyp}$ The staff were very friendly and helpful. The room was clean and comfortable.
@@
The central idea of this study is to inspect a hypothesis output, generated with a default decoding strategy, with stochastic samples obtained under varying temperature. At low temperature settings, we observe that the sampled outputs are highly similar to the hypothesis if the model is not hallucinating, and moderately dissimilar if it is. We believe this occurs because of the following reason: when a clear optimal candidate token exists, the default decoding strategy and stochastic sampling tend to favor the most probable token. However, in cases where no single candidate clearly dominates (often indicative of hallucination), the outputs are more likely to diverge. In contrast, at high temperature, generation becomes increasingly random, and sampled outputs tend to be substantially different from the default output regardless of the model’s behavior. This behavior is illustrated in Table 1. As a consequence, plotting the similarity between the default output and the stochastic outputs yields two distinct curves: either a monotonously increasing trajectory in the non-hallucinating scenario, or a $U$-shape curve in the hallucinating scenario. This can be visualized in Figure 1.
Figure 1: MMD-Flagger at work on the LFAN-HALL dataset: two outputs, generating two MMD trajectories. In orange, the trajectory for the hallucinated translation (see Table 1), in maroon, the trajectory corresponding to the correct one. Text boxes show examples of $y_{\rm hyp}$ and samples from $Y_{\rm sto}^{\tau}$ (two of 25 samples), generated by stochastic decoding with temperatures $\tau$. We flag $y_{\rm hyp}$ as a hallucination when the minimum value of $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ (the vertical black dashed line) falls within the blue region. Otherwise, the output is deemed plausible (minimum falling into the blue region).
Drawing on these findings, we introduce a novel hallucination detection method: MMD-Flagger. The key idea of the method is to simply i) plot the similarity between default output and stochastic samples as a function of the temperature, and ii) automatically detect whether this curve is monotonously increasing or $U$-shaped. In the latter case, we flag the output as hallucination. As a measure of similarity, we propose to use the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012). As a shape detection mechanism, we simply estimate the inflection point of the trajectory and check whether it is located near 0 (see Figure 1).
The paper is organized as follows. In Section 2, we detail the operating procedure of MMD-Flagger after recalling the necessary background on temperature sampling and MMD. We benchmark our method against several natural competitors in Section 3, before concluding in Section 5. The code for all experiments is provided in code.tar.gz and will be publicly available after publication.
@@ 1.1 Related work
Hallucination detection in LLMs has attracted significant attention, with a variety of methods proposed in recent years. We follow the categorization outlined by Park et al. (2025) and highlight the approaches most relevant to our work.
Consistency-based methods. These methods detect hallucinations by comparing multiple generations from the same prompt under different decoding strategies. Our proposed method, MMD-Flagger, belongs to this category. MC-DSIM (Fomicheva et al., 2020; Zerva et al., 2021) leverages Monte-Carlo dropout (Gal and Ghahramani, 2016) to generate multiple outputs and then computes an uncertainty score using a similarity metric in natural language space. Semantic Uncertainty Nikitin et al. (2024) similarly collects stochastically sampled outcomes, builds a graph representation, and derives an uncertainty score from graph structure. SelfCheckGPT (Manakul et al., 2023) compares candidate generations against stochastically sampled alternatives. SelfCheckGPT and Semantic Uncertainty require external models (e.g. inference models) for building a graph or scoring. In contrast, MMD-Flagger operates without external resources.
Logits-based methods. These approaches use log-probabilities to assess the fluency of the generated token sequence (Ren et al., 2023). However, hallucinations that are syntactically plausible can still receive high log-probabilities, making them difficult to detect. In addition, setting thresholds for logits-based scores presents practical challenges.
Hidden-State-based methods. Several works exploit internal representations to identify hallucinations. Cross-attention patterns in encoder–decoder models have been used for detection (Guerreiro et al., 2023a; Ferrando et al., 2022; Dale et al., 2023), though such methods are not directly applicable to decoder-only architectures. Other approaches apply matrix decomposition techniques such as SVD to hidden states (Chen et al., 2024; Du et al., 2024; Park et al., 2025), isolating subspaces thought to capture hallucination-specific features. These methods assume hallucinations appear in limited proportion within the dataset; if hallucinations dominate or are absent, the identified subspace may fail to characterize them effectively.
Probing-based methods. Recent work probes hidden states directly to capture differences between grounded and hallucinated responses (Oblovatny et al., 2025; Bouchekir et al., 2025; Xu et al., 2023). For instance, Oblovatny et al. (2025) employ MMD to measure distances between prompt and response representations, assuming that grounded responses exhibit larger deviations due to deeper reasoning, whereas hallucinations arise from superficial rephrasing with smaller deviations. However, this assumption is limiting: fabricated responses can also diverge significantly from the prompt, leading to false positives.
@@ 2 Methodology
In this section, we describe the proposed method, MMD-Flagger. Given an input document $x$, the starting point of the method is to generate the hypothesis output $y={\rm LLM}(x)$ to be inspected. The next step is to generate additional documents by stochastic sampling at various temperature parameters, which is detailed in Section 2.1. Then, the MMD between $y$ and the newly created groups is computed for each temperature parameter. We detail this computation in Section 2.2. At this point, for a given input / output pair $(x,y)$, we have obtained a series of MMD values corresponding to different temperature settings. We explain our procedure to inspect this trajectory in Section 2.3.
2.1 Decoding strategies and stochastic sampling
In this section, we present the generation process by LLMs considered in this work and detail the role of the decoding strategies.
Text generation with LLMs. We consider LLMs based on the Transformer architecture (Vaswani et al., 2017). Given an input sequence of tokens $x$, the model first processes the sequence through multiple layers of self-attention and feed-forward networks. In encoder–decoder models (Sutskever et al., 2014), the encoder maps $x$ into a contextual representation, which is then attended to by the decoder. In decoder-only models, the same architecture autoregressively encodes and generates tokens within a single stack of Transformer blocks, where causal masking ensures that predictions for each position depend only on preceding tokens.
The decoder generates an outcome sequence autoregressively, conditioned on the prompt and the previously generated tokens. Formally, each token is sampled as
$Y_{t}\sim\Phi\left(Y_{t-1},\ldots,Y_{1},\mathcal{H}(x)\right),$
where $\mathcal{H}(x)$ denotes the model’s hidden representation of the prompt, and $\Phi$ is a stochastic sampler applied to the model’s predictive distribution. In case of the encoder-decoder models, $\mathcal{H}(x)$ is the encoder output.
At each step, the model produces a vector of logits in $\mathbb{R}^{V}$, where $V$ is the vocabulary size. These logits are transformed by a softmax layer into a probability distribution $P\in\mathbb{R}^{V}$ over the next token, from which the sampler $\Phi$ selects the output. This autoregressive process continues until an end-of-sequence (EOS) token is generated.
@@ 2.1 Decoding strategies and stochastic sampling
Stochastic sampling with temperature. The stochastic sampling strategy collects multiple outputs for a given input $x$ that are used to inspect with the hypothesis output $y$. That is, sample a random variable with distribution given by $P$ and select the associated token. We denote $Y_{\text{sto}}=\text{LLM}_{\text{sto}}(x)$ the random output obtained by stochastic sampling.
@@
A variation is introduced by a temperature parameter $\tau>0$. The softmax transformation is modified into $P_{i}=\mathrm{exp}(z_{i}/\tau)/\sum_{j=1}^{V}\mathrm{exp}(z_{j}/\tau)$, where $z_{i}$ represents the logit score for the $i$-th token. As recalled in Section 1, high temperature gives a lot of variability (Yehuda et al., 2024) (equivalent to uniform sampling in the limit), while low temperature degrades to argmax sampling.
Decoding strategies for the hypothesis. We introduce the decoding strategies for selecting the token associated with the highest $P_{i}$ as output. These decoding strategies are used to obtain a token sequence that is inspected if the hallucination exists. We denote this generation process as $y_{\rm hyp}={\rm LLM_{\rm hyp}}(x)$.
A systematic study (Shi et al., 2024) shows that no universally optimal decoding strategy exists; the best choice depends on the model architecture, alignment process, computational resources, and application requirements. As a general guideline, beam search is recommended for aligned models, while stochastic sampling strategies are preferred when computational resources allow, as they often improve diversity and robustness. Beam search is a deterministic decoding algorithm that maintains a set of $w$ candidate sequences, updating them at each time step by retaining the $w$-highest probability continuations (Meister et al., 2020). Although deterministic, beam search and its variants are plausible inspection targets because the token candidates are fundamentally selected based on probability distributions. This probabilistic foundation ensures that MMD-Flagger can inspect hypothesis outputs generated by any decoding method, including deterministic approaches like beam search as well as stochastic sampling strategies such as Top-$k$ sampling (Fan et al., 2018) and nucleus sampling (Top-$p$) (Holtzman et al., 2020).